Run plan history CREX

From PREX Wiki
Jump to: navigation, search



CREX Main << Run Coordinator Tools << RunPlanDaily_CREX

Copy weekly run plan tables here.


Day, Date Shift Tasks (Responsible Party) Status
Monday, November 25 Owl No shift
Day No shift
Swing No shift
Tuesday, November 26 Owl No shift
Day Compton work
Swing Compton work
Wedsday, November 27 Owl No shift
Day No shift
Swing No shift
Thursday, November 28 Owl No shift
Day Eat Turkey low current commissioning
Swing Eat more...
Friday, November 29 Owl No shift
Day No shift
Swing
Saturday, November 30 Owl No shift
Day No shift
Swing No shift
Sunday, December 1 Owl No shift low current commissioning
Day Moller Measurement
Swing Moller Measurement
Monday, December 2 Owl No shift
Day Magnet work
Swing Magnet work
Tuesday, December 3 Owl No shift
Day Magnet work
Swing Magnet work
Wednesday, December 4 Owl No shift
Day Magnet work...
Swing Magnet work...until Jack ready to leave 4-8pm. Then 5pass tuning into A.
Thursday, December 5 Owl 5pass tuning into A.
Day RF stress test to A/RF fixing Burned up 3 raster power supplies (ended test at 4:15pm!)
Swing Go to 1 pass; Septum problem (only up to 400 A); collimator centering; polarity check Set HRS tune P for this septum setting for polarity check--not conclusive
Friday, December 6 Owl spectrometer setup did not allow target alignment checks Beam gone at 3:40am due to Network issue
Day Hall surveyed and in controlled access all day; remove diffusor at end? Septum control software fix; compton e-det work; network issues run down; raster sync problem fixed
Swing lock up hall; reverse septum polarity; raster size and hole locations Left arm NMR probe lock not working--will have to use Hall probe; couldn't put any beam on Ca-40 due to dump FSDs (and beam optics problem -- MOLLER quads set too high)
Saturday, December 7 Owl Tried ion-chamber calibration, but failed; started spec commissioning First pass optics DB looks good; started septum setting and inner edge searches. Right dipole tripped so need to redo inner edge search
Day Jay re-establishes Yves nominal crex beam transport setup; re-check raster size (1.8 x 1.4 now is ~2 mm x 2 mm); new hole centers (did not change much); ion-chamber calibrations done for No target, C-1%, Ca-40, and Pb#4 (on cold ladder); resumed spec. commissioning Found out reason for raster optics issues: Moller quads not set correctly after 5-pass backout causing wide and short raster settings that evolve to tall and narrow raster at the dump (but 2x2 at the target)...
Swing Re-did inner-edge septum scan, spot minimization chase (Sangwa, Ciprian) with quartz checkout (Devi); (call Bob) start optics commissioning We established the preliminary 2mm x 2mm (+/-0.1mm) raster (1.8 x 1.4 MCC), but should be double checked/tweaked before going high current on Calcium
Sunday, December 8 Owl Continued Optics commissioning: dp scans with water-cell measurements most of shift Experts coming in: Bob, Chandan, and Siyu
Day Authorize 150uA; finish optics commissioning; Followup optics studies performed: Q1 scans (-10 to -20%) at +5% septum, nominal tune "P", find best Q1 setting, then scan Q3's to effectively minimumize the FP y widths The established tune "P" for crex gave large x/th coupling, a lack of the "bow-tie" FP theta v. phi shape, and anomalous structures in the dispersive-x FD distributions; seems to be some quad strength in the septum field that is cancelled out by lowering Q1
Swing Established CREX tune: "tune P" with +5% septum, -17%/-15% LQ1/RQ1, -5% Q3s. Redo dp optics scan but without w-cell runs, y-target reconstruction Right dipole tripped OFF at 10:15pm, started ramping back up at 3AM. Aborted sieve-in optics.
Monday, December 9 Owl Took sieves out and proceeded with frame finding/quartz alignment and rate activities. Bob is working to fix left arm spot. R dipole back and stable by 6am, however N Lical went down at 5:30am due to vacuum issue(s) Experts coming in: Bob, Chandan, and Devi
Day SL shift cancelled; TO still required. Hall surveyed and in Controlled Access for opportunistic work Compton work in the hall; Accelerator down at least until 8 AM next day and may restore beam delivery by 12pm tomorrow
Swing SL shift cancelled; TO still required
Tuesday, December 10 Owl SL shift cancelled; TO still required
Day Still investigating vacuum issues in the N. Linac. Spectrometers up and running smooth Hall in Controlled Access for more Compton e-det and laser work; Hall is ready for beam. Caryn is coordinating a solution to the raster size and FSD lower limit issues
Swing N. Linac problem fixed; waiting for beam Around 10pm, accel was ready but Ops-controlled valves in the A-line downstream of the wall were not remotely controlable and no remote status update. Problem was investigated, but not properly escalated and continued into Owl
Wednesday, December 11 Owl No beam all shift; commincation b/t MCC, Hall, and techs was insufficient regarding this valve problem...Shifters could not get any info on the problem or its progress towards resolution... Continued confusion about what to do with valve problem...Pete Francis get's in at 4am and starts diagnosing
Day No beam all shift; trouble shooting remote communication problems on beamline valves and vacuum pressure monitors. Pete determined it's not a networking issue (other systems on the crate work); it's localized to the status readback and controller cards specific for the valves in the CAMAC crate. Isolated problem to outdated software libraries that crept into system during the RF stress test raster fixes. This problem also affected all downstream of target valves--they were reported as open, but where closed... There is also a problem with 3 cold-cathode probes--not reading back properly (too good and quiet).
Swing When beam is back: Continue spectrometer commissioning with low current (steps 11-13): Frame finding/centering, Q2 and quartz rate study on C-1%, Ca-40, and Ca-48 (Devi or Chandan); verify quartz alignment for Ca-48 (Devi); A_T "box" scans (Devi) Expect very similar rates for all three targets: ~700k Hz per uA T1 scalar rate
Thursday, December 12 Owl Finish low current plans, if necessary; then begin a modified high current checkout (note that the ion chamber calibration has already been completed); proceed with steps 1) Basic setup, except with 1% C target and not steering through Compton 2) UNSER calibration (also with 1% C target); then thermal break-in procedure for targets; if these steps are complete continue taking data on Ca-48 (carefully monitoring temps and widths) What other beam instrumentation could be checked-out during thermal break-ins?
Day Spin dance start at 7am if Hall C ready Notified by Dave that Spin Dance will likely be moved to Friday
Swing if no spin dance during Day, then continue high-current commissioning When Compton laser system is ready OR we are ready for production data: Preferred order is -Tune through the Compton chicane, ion chamber (re-)calibration, FFB, and BMW target out--phase advance setup and for production target.
Friday, December 13 Owl high-current commissioning
Day Spin dance cancelled; production when have beam
Swing high-current commissioning/pre-production Goal is to get the beamline instrumentation ready for PQB to get a look at regression and widths, etc. (pre-production)
Saturday, December 14 Owl Production
Day Production, change IHWP, check target pressure with large raster
Swing Production


Run plan history starting January 2020

-
Day, Date Shift Tasks (Responsible Party) Status
Monday, Jan 6 Owl
Day
Swing Moller Lock, FFB low current, Optics Raster Iy issue resolved, all same as Dec.
Tuesday, Jan 7 Owl Optics Larm det ul base replaced
Day Mott, Spin Dance Vac gauge readout 1-e12, vac adjacent fine (not urgent, to be fix later)
Swing Spin Dance, 150uA to dump Experts: (Simona, Eric) , (Caryn, Devi)
Wednesday, Jan 8 Owl 150uA, bcm/det calibration, damping measure, bmw initialization...production Experts: Caryn, Devi, Ye
Day Hall Access, fix compton T controller. OPS down:RF recovery, NL RF cav gradient, energy loss, broken klystron, failed arc det.
Swing Tune through chicane attempt, bleed-through from B discovered, returned to straight through end of shift, but compton rates too high Experts: Compton(Dave, Adam, Amali), BMW(Ye,Victoria,Caryn)
Thursday, Jan 9 Owl Low current BMW tune 1H01 test Couldn't send beam due to compton rates
Day halld laser inj access, bleedthrough B into A injector work prebuncher, chopper.
Swing compton tuning through chicane...but match into hall was wrong. Couldn't send beam due to spot size limitations.
Friday, Jan 10 Owl Tuning match into Hall
Day "Production", injector adjustment, injector match
Swing injector match, hall match, hall on design, lower compton rates straight through, compton chicane tuning attempt ran out of time
Sat Jan 11 Owl Production, FFB gain issues (hi search limit), yield det decrease and width increase over night
Day detector alignment check and left dipole current adjustment. Production. LQ1 trip, recycle.
Swing 7:53PM HA to power permit, beam -> C
Sun Jan 12 Owl (figure out HC feedback again)
Day Radcon survey, Compton Laser work (TO HC feedback )
Swing Compton Laser work (TO HC feedback )
Mon Jan 13 Owl (TO HC feedback )
Day Vacuum group work ion pump, Compton laser work (TO HC feedback )
Swing Compton laser work, Hall Lockup, tell OPS PQB checkout in INJ when HA coming back online (TO HC feedback )
Tues Jan 14 Owl Production,
Day Compton chicane tuning, commissioning, Production Don's Run Plan
Swing BMW fine tune, Production Don's Run Plan
Wed Jan 15 Owl Production,
Day Compton chicane tuning, commissioning, Production Don's Run Plan
Swing BMW fine tune, Production Don's Run Plan
Th Jan 16 Owl Production,
Day Hall C tuning, production
Swing Production,
Fr Jan 17 Owl Compton setup, Production,
Day Machine tuning; return to production
Swing Production
Sat Jan 18 Owl Production,
Day Production
Swing Production
Sun Jan 19 Owl Down due to target issue, controlled access
Day Down due to target issue, controlled access
Swing Down due to target issue, controlled access
Mon Jan 20 Owl Down due to target issue, controlled access
Day Down due to target issue, controlled access
Swing Down due to target issue, controlled access
Tue Jan 21 Owl Down due to target issue, controlled access
Day Down due to target issue, controlled access
Swing Down due to target issue, controlled access
Wed Jan 22 Owl Down due to target issue, controlled access
Day Down due to target issue, controlled access Ca40, Ca48 target reinstalled
Swing
Thu Jan 23 Owl
Day Target cooldown
Swing Beam restoration, target commissioning, detector alignment check, Qsq measurement
Fri Jan 24 Owl BMW check, target thermal commissioning
Day Production!
Swing Production!
Sat Jan 25 Owl
Day
Swing
Sun Jan 26 Owl
Day
Swing
Mon Jan 27 Owl
Day Moller measurement
Swing Moller measurement
Tue Jan 28 Owl
Day
Swing
Wed Jan 29 Owl
Day
Swing
Thu Jan 30 Owl
Day
Swing
Fri Jan 31 Owl
Day production
Swing production
Sat Feb 1 Owl production
Day production
Swing production
Sun Feb 2 Owl production
Day Moller measurment
Swing Moller measurment
Mon Feb 3 Owl production
Day (Hall C spin dance) (Hall C spin dance)
Swing
Tue Feb 4 Owl
Day
Swing
Wed Feb 5 Owl
Day
Swing Moller was scheduled, but postponed

Restored beam all shift (BLA calibration, Ion Chamber calibrations)

Thu Feb 6 Owl Production
Day Production
Swing Production (until 8 PM)

Moller (starting at 8 PM)

Network issues caused lost Acc IOC communications

(Hall A experienced several brief but quickly resolved connection errors as well)

Fri Feb 7 Owl (Moller) No beam - recovering from Acc IOC communications problems
Day (Production) Backed off Moller setup (Moller postponed)

Restore beam to hall Develop and test new Compton lock

Swing (Production) Hardly any production running

Added new Hall A Target lock behavior

Saturday Feb 8 Owl (Production) No beam - more Accelerator IOC unresolved issues
Day Moller Moller setup started mid shift
Swing Moller
Sunday Feb 9 Owl Production
Day Production
Swing Production
Monday Feb 10 Owl Production
Day Setup vertical polarization

Transverse

Swing Transverse
Tuesday Feb 11 Owl Transverse
Day Transverse
Swing Transverse
Wednesday Feb 12 Owl Transverse
Day No beam (Hall B at 1 pass)
Swing No beam (Hall B at 1 pass)
Thu Feb 13 Owl No beam (Hall B at 1 pass)
Day No beam (Hall B at 1 pass)
Swing No beam (Hall B at 1 pass)
Fri Feb 14 Owl No beam (Hall B at 1 pass)
Day No beam (Hall B at 1 pass)
Swing No beam (Hall B at 1 pass)
Sat Feb 15 Owl No beam (Hall B at 1 pass)
Day No beam (Hall B at 1 pass)
Swing Pass change and tuning beam to Hall A
Sun Feb 16 Owl Production
Day Production
Swing Production
Mon Feb 17 Owl Production
Day Set up Wien Flip Left
Swing Production
Tue Feb 18 Owl Production
Day Production
Swing Production
Wed Feb 19 Owl Production
Day Production
Swing Production